Calculate the mole fraction of naoh in a solution containing 20 g of naoh per 360 gram of water
1
Expert's answer
2020-07-17T05:40:48-0400

Mole fraction refers to the ratio of moles of NaOH and number of moles of solution

Hence, mole fraction = (moles of NaOH)/(moles of solution)


Number of moles = mass of solute/molar mass of solute

mass of NaOH = 20g, and molar mass = 40g

moles of NaOH = (20/40) = 0.5

moles of solution = moles of NaOH + moles of water

moles of water = (360/18) = 20

mole fraction = 0.5/(0.5 + 20) = 0.024 (expressed to 3 decimal places)